OpenStudy (anonymous):

Peanuts worth $1.50 per pound were mixed w/ cashews worth $2.25 a pound. How many pounds of each were used if there're 20 pounds at $1.80 per pound?

OpenStudy (anonymous):

x=pounds of peanuts y=pounds of cashews we set up a system of equations: 1.50x+2.25y=20(1.80) and x+y=20 1.5x+2.25y=36 we can set x to equal 20-y and substitute 1.5(20-y)+2.25y=36 30-1.5y+2.25y=36 .75y=6 y=8 if y=8 then x=12, so there are 12 pounds of peanuts and 8 pounds of cashews